how do you solve this equation 6r-1+6r=11?

Answer by ReadingBoosters(3246)

Starting with
6r - 1 + 6r = 11
Combine like terms
6r + 6r = 11 + 1
12r = 12
Divide both sides by 12
r = 1
.........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.
www.MyHomeworkAnswers.com